The complex function can be modelled by equation y=3x^2 x e^x3 Find the indefinite integral of function using substitution method
1
Expert's answer
2020-03-05T11:42:45-0500

The function can be modelled by the equation given by

"y = 3x^2 e^{x^3}"

The Indefinite integral of this equation is:

"\int y \, dx"

"= \int 3x^2 e^{x^3} \, dx"

Let us assume that:

"x^3 = u \\
\Rightarrow 3x^2 \, dx = du"

Substituting "x^3 = u" into the integration and integrating with respect to "du" , we have:

"= \int e^u \, du \\"

"= e^u + C \hspace{1 cm}" [ "\because \int e^x \, dx = e^x + C \,\," (where C is a constant of integration) ]

Undo substitution and we have:

"= e^{x^3} + C \hspace{ 1 cm} \left[ \because u = x^3 \right]"